# Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries

### v1

@@ 7 Conclusion
In this paper, we present Visual-RAG, a challenging VQA benchmark that requires visual knowledge as evidence to answer questions. Different from previous works that include images as part of queries and let models retrieve textual knowledge, our Visual-RAG is formed with text-only queries and enforces text-to-image retrieval for challenging visual knowledge. We evaluate 8 popular MLLMs under various multimodal RAG settings. The results indicate that images can serve as knowledge sources for answering visual knowledge intensive questions, but the models struggle to effectively extract visual knowledge from a single clue image. Surprisingly, the current models are able to reason across clue and non-clue images for better understanding of the questions. As one of the pioneer works in visual-knowledge centric multimodal RAG, our benchmark can help promote the research and development of MLLMs in extracting fine-grained visual knowledge and enhance cross-modal retrieval in a more realistic setting.
Limitations
While our Visual-RAG addresses the lack of visual knowledge intensive QA benchmarks, it has several design constraints. First, our query generation and manual rewriting process focus on single-hop questions only. Although a small subset of queries might involve simple comparisons within the same image, such as checking the colour of a bird’s upper beak vs. lower beak, there are no truly multi-hop queries requiring reasoning across multiple images. Designing such multi-image or even cross-species queries would demand significantly more expert knowledge, and we leave this as an avenue for future work.
Secondly, Visual-RAG focuses exclusively on organisms, which raises questions about the broader generalizability of our findings. We chose the domain of organisms because they allow for factual queries regarding visual features that remain consistent across multiple individuals of a species, yet are present in only a minority of images. Other commonly used domains in knowledge-intensive VQA such as buildings, artworks, news event, etc., do not always meet these criteria, as answers may not hold universally for all instances or there exists only one instance, and the required visual features are often too prevalent or too trivial to foster a genuinely challenging retrieval scenario.
Lastly, the overall scale of Visual-RAG is limited compared to larger benchmarks like InfoSeek and Encyclopedic-VQA. Their heuristic-driven methods for automatically generating queries and answers do not readily apply to our specialized setting, which hinges on expert-informed checks to ensure the presence (or absence) of specific visual features. Despite these constraints, Visual-RAG represents an important step toward truly visual knowledge intensive QA, and we hope it will inspire further research into more complex, multi-hop, and cross-domain tasks.
Ethical Considerations
@@
Appendix A Relation with previous works
@@
Knowledge intensive VQA benchmarks focusing textual knowledge, including but not limited to: ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (Mensink et al., 2023), are discussed in related works. These benchmarks contain questions paired with images of certain entities, and knowledge for such entities can be found in an external textual knowledge base (e.g., Wikipedia). The questions are deliberately constructed to omit the entity names (e.g., “How many feet tall does the plant grow to?” instead of “How many feet tall does Acacia paradoxa grow to?”).
@@
Typically, given a question and its associated image, a model is expected to: 1) identify the entity depicted in the image, 2) retrieve relevant documents or passages from a knowledge base that pertain to both the question and the entity, and 3) generate an answer based on the question, image, and retrieved documents. Step 1) and 2) is usually carried out by leveraging vision-language pretrained encoders, such as CLIP (Radford et al., 2021), to facilitate cross-modal (image-to-text) retrieval.
@@
While the above mentioned benchmarks are all claimed to be Knowledge-Intensive VQA benchmarks, it is noteworthy that the knowledge to be retrieved is textual knowledge. The image is only served as an anchor of the entity. While cross-modal retrieval was commonly used, the knowledge retrieval process can actually be achieved in uni-modal manner. A recent work (Yan and Xie, 2024) achieved “SoTA” performance on InfoSeek and E-VQA employing image-to-image retrieval within a knowledge corpus comprising Wikipedia pages and their associated images. Entity recognition is performed by identifying the most visually similar Wikipedia image to the query image, which automatically enables retrieval of the related knowledge document (i.e., the Wikipedia page containing that image). Answers are then generated using a text-only LLM, augmented with the relevant paragraph from the Wikipedia page, which is extracted through fine-grained reranking.
@@
WebQA dataset (Chang et al., 2022) included questions that require retrieving images and utilizing visual knowledge to generate answers. However, its retrieval process is heavily reliant on image captions, as the images themselves typically lack indicators identifying them as clue images for answering a question. For instance, considering the question: “Are the land dinosaurs guarded by rail in both the Display Museum of Natural History in University of Michigan and the Museo Jurassic de Asturias?” Without captions, the model would be unable to discern which dinosaur fossil image corresponds to which museum.
Table 5 shows the retrieval scores in our implementation of retrieval experiments on WebQA benchmark for queries having images as knowledge source. While the queries are text-only, aligning with our setting, the text-to-image retrieval scores significantly lag behind text-to-caption retrieval, especially for lower $k$.
@@
Text-to-Image 15.17 37.08 47.63 58.43 72.32
Text-to-Caption 43.25 68.74 76.86 83.79 89.92
Table 5: Retrieval scores for WebQA dataset.
MRAG-Bench (Hu et al., 2024) shares a similar objective to ours by proposing a benchmark that requires retrieving visual knowledge to solve questions. However, despite this similarity, the focus remains on entity recognition. Most tasks are structured such that an image displaying an incomplete or partially visible entity is provided, and the retrieval system is tasked with identifying images that depicts the entire entity. This approach primarily enhances the MLLM’s ability to recognize entities rather than addressing broader knowledge-intensive tasks. Furthermore, non-clue images in MRAG-Bench are from different entities, which are visually distinct, in contrast to the numerous hard negatives in our Visual-RAG .
Concurrent to our work, the iNaturalist team released 2024 version of their natural image dataset, iNat24 (Vendrow et al., 2024), coming with a text-to-image retrieval benchmark named INQUIRE, containing 250 queries. Though, the queries in INQUIRE are designated for image retrieval only, and are not QA style questions that have explicit answers. As iNat24 and iNat21 share the same taxonomy classes of organisms, we do consider adding the new images in iNat24 to our benchmark.
We also note a related line of research (Tanaka et al., 2023; Van Landeghem et al., 2023) on extracting knowledge from document images (e.g., PDF files, Powerpoint slides), conventionally referred to as OCR-based VQA. Although these tasks also require models to retrieve relevant page and interpret images to answer textual queries, the underlying knowledge is primarily textual—obtained via OCR. This focus diverges considerably from our goal of visual feature extraction, where visual cues are inherently sparser in the wild.
Appendix B Construction of Minimum Textual Knowledge Base
@@
We then applied basic text chunking to these articles, splitting them into chunks of approximately 150 tokens. Each chunk is bounded within a single section of the original article to preserve contextual integrity. Sections such as “Reference”, “See also”, “External links”, etc., which lack informative content, were excluded. The chunks are cut upon period mark, so that they won’t end or start with incomplete sentence. Two adjacent sections, having total length less than 150 tokens will be merged, with the section titles kept.
@@ Appendix C Annotation details
For the query generation, we mentioned that we have gone through a human annotator originated refined filtering and rewriting process. The prompt template is shown in Table LABEL:tab:prompt_qgen, and following shows the example prototype queries generated with the sample prompt:
@@
Figure 4: Image of Bromus catharticus flower anther. Owned by original uploader at www.inaturalist.org, shared under CC-BY-NC license.
@@
After query generation and rewriting, and coarse-level filtering described in section 4, we dispatch the validated queries to the next round of human annotation of image labels. Figure 5 shows the annotation interface for image labelling. [^8] Similar interface is used for query filtering and rewriting annotator to browse, except the image matrix is more dense and displaying only thumbnail sized images. Annotators are required to only label “Y”, to indicate that this image contain the queried feature and is a clue to the query. Leaving the checkbox blank will be considered as negative label. This design aims to save annotators’ effort, as clue images are assumed to be the minority. The rough clue image rate predicted by MLLM in the previous step is also provided (“Y_Rate: 0.003” after the question text) as reference for annotators to estimate the query difficulty. In the meanwhile, annotators provide query answers upon finding a clue image, as shown in Figure 6. Annotators are allowed to skip for a query when finding more than 9 clues after the first 27 images, indicating this query likly fail the clue image threshold. The order of images are randomly shuffled.
Figure 5: Annotation interface for image labelling
Figure 6: Annotation interface for image label and query answer
Appendix D Prompt Template
This section list the prompt templates used through out our experiments.
@@ Appendix D Prompt Template
A likely related paragraph from Wikipedia is provided for you to better understand the organism. Please avoid question regarding visual features explicitly mentioned in the Wikipedia paragraphs. (e.g., paragraph: “The seeds are colored in red” -> Don’t ask “What color is the seed”; paragraph: “The plant have symbolic seed pods” (but doesn’t mention its color) -> You may ask “What shape/color is the seed pod”)
@@
1. What is the common color of tip of leaf of wall-screw moss (scientific name: Tortula muralis)? (Explanation: The moss can be distinguished from similar mosses by its ḧair-pointedl̈eaves.) (Feature: color of leaf tip)
@@
2. What pattern can be observed on the tail feather of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) from dorsal (back) view? (Feature: pattern on tail feather from dorsal view)
More ICL Examples omitted
@@
Wiki Paragraph:
Bromus catharticus is a species of brome grass known by the common names rescuegrass, grazing brome, prairie grass, and Schrader’s bromegrass. The specific epithet catharticus is Latin, meaning cathartic. The common name rescuegrass refers to the ability of the grass to provide forage after harsh droughts or severe winters. The grass has a diploid number of 42 …… Omitted due to space. Actual prompt includes whole description passage.
@@
Table 6: Prompt Template for Query Generation. Texts in italic and purple are not part of prompt.
@@
<image> <image> <image>…
@@
Please evaluate the answer to a question, score from 0 to 1. The reference answer is provided, and the reference is usually short phrases or a single keyword. If the student answer is containing the keywords or similar expressions (including similar color), without any additional guessed information, it is full correct. If the student answer have missed some important part in the reference answer, please assign partial score. Usually, when there are 2 key features and only 1 is being answered, assign 0.5 score; if there are more than 2 key features, adjust partial score by ratio of correctly answered key feature. The reference answer can be in the form of a Python list, in this case, any one of the list item is correct.
If student answer contain irrelevant information not related to question, mark it with “Redundant”, but it does not affect score if related part are correct. (e.g. Question: what shape is leave of Sanguinaria canadensis, Student Answer: shape is xxx, color is yyy, this is Redundant answer)
If student answer contain features not listed in reference answer, mark it with “Likely Hallucination” and deduct 0.5 score. (e.g., Reference Answer: black and white. Student Answer: black white, with yellow dots, “yellow dots” is not mentioned in reference)
Separate the remarks with score using “|”, that is, use the syntax of: “Score: score | Likely Hallucination”, “Score: {score}”, “Score: {score} | Likely Hallucination | Redundant”, etc. If any explanation on why giving the score is needed, do not start a new line and append after remark with brackets, e.g. “Score: {score} | Redundant | (Explanation: abc)”.
@@
Reference Answer: black eye-round feather, white stripe above eyes, sometimes connected to the white forehead
@@
Question: …
More ICL Examples omitted
@@
Table 8: Prompt Template For Automatic Evaluation. Note that for evaluation, no images is provided, and it becomes pure-text open-ended QA evaluation. Texts in italic and purple are not part of prompt.

### v2

@@ 6 Conclusion
In this paper, we present Visual-RAG, a visual knowledge-intensive VQA benchmark that requires models extracting evidence encoded in images to answer questions. Different from previous works that include images as part of queries and let models retrieve textual knowledge, our Visual-RAG is formed with text-only queries and enforces text-to-image retrieval for challenging visual knowledge grounded question answering. We evaluate eight popular MLLMs under various multimodal RAG settings. The results indicate that images can serve as knowledge sources for answering visual knowledge-intensive questions, but the models struggle to effectively extract visual knowledge from noisy retrieved contexts with hard negatives introduced. We further quantify model behavior with a unified efficiency metric $\mathrm{gCUE}$, measuring models’ robustness at different numbers of images. As one of the pioneer works in visual-knowledge centric multimodal RAG, our benchmark can help promote the research and development of MLLMs in extracting fine-grained visual knowledge and enhance cross-modal retrieval in a more realistic RAG setting.
Ethical Statement
@@
Appendix A Extended Comparison with Related Works
@@
Knowledge intensive VQA benchmarks focusing textual knowledge, including but not limited to: ViQuAE (Lerner et al. 2022), InfoSeek (Chen et al. 2023) and Encyclopedic-VQA (Mensink et al. 2023), are discussed in related works. These benchmarks contain questions paired with images of certain entities, and knowledge for such entities can be found in an external textual knowledge base (e.g., Wikipedia). The questions are deliberately constructed to omit the entity names (e.g., “How many feet tall does the plant grow to?” instead of “How many feet tall does Acacia paradoxa grow to?”).
@@
Typically, given a question and its associated image, a model is expected to: 1) identify the entity depicted in the image, 2) retrieve relevant documents or passages from a knowledge base that pertain to both the question and the entity, and 3) generate an answer based on the question, image, and retrieved documents. Step 1) and 2) is usually carried out by leveraging vision-language pretrained encoders, such as CLIP (Radford et al. 2021), to facilitate cross-modal (image-to-text) retrieval.
@@
While the above mentioned benchmarks are all claimed to be knowledge-intensive VQA benchmarks, it is noteworthy that the knowledge to be retrieved is textual knowledge. The image is only served as an anchor of the entity. While cross-modal retrieval was commonly used, the knowledge retrieval process can actually be achieved in uni-modal manner. Yan and Xie (2024) achieved “SoTA” performance on InfoSeek and E-VQA employing image-to-image retrieval within a knowledge corpus comprising Wikipedia pages and their associated images. Entity recognition is performed by identifying the most visually similar Wikipedia image to the query image, which automatically enables retrieval of the related knowledge document (i.e., the Wikipedia page containing that image). Answers are then generated using a text-only LLM, augmented with the relevant paragraph from the Wikipedia page, which is extracted through fine-grained reranking.
@@
WebQA dataset (Chang et al. 2022) included questions that require retrieving images and utilizing visual knowledge to generate answers. However, its retrieval process is heavily reliant on image captions, as the images themselves typically lack indicators identifying them as clue images for answering a question. For instance, considering the question: “Are the land dinosaurs guarded by rail in both the Display Museum of Natural History in University of Michigan and the Museo Jurassic de Asturias?” Without captions, the model would be unable to discern which dinosaur fossil image corresponds to which museum. Table 4 shows the retrieval scores in our implementation of retrieval experiments on WebQA benchmark for queries having images as knowledge source, evaluated with clip-vit-large-patch14-336. While the queries are text-only, aligning with our setting, the text-to-image retrieval scores significantly lag behind text-to-caption retrieval, especially for lower $k$.
@@
T2I 15.17 37.08 47.63 58.43 72.32
T2T 43.25 68.74 76.86 83.79 89.92
Table 4: Retrieval scores for WebQA dataset. T2I: Query-to-Image retrieval. T2T: Query-to-Caption retrieval. T2T retrieval significantly outperform T2I, indicating crucial identifier for images being the clue reside in caption.
In addition to MRAG-Bench introduced in the main paper, another recent work M2RAG (Liu et al. 2025) gathers data samples from existing benchmarks, namely WebQA introduced in previous paragraph and Factify (Mishra et al. 2022), a multimodal fact verification dataset. The multi-modal QA subtask introduced partially overlap with ours, while the queries are repackaged instances drawn from the original WebQA dataset.
The iNaturalist team released 2024 version of their natural image dataset, iNat24 (Vendrow et al. 2024), coming with a text-to-image retrieval benchmark named INQUIRE, containing 250 queries. Though, the queries in INQUIRE are designated for image retrieval only instead of QA style questions that have explicit answers. As iNat24 and iNat21 share the same taxonomy classes of organisms and contain newly uploaded images not included in iNat21, we do consider adding the new images in iNat24 to our benchmark. We conduct preliminary experiment using iNat24 images as retrieval corpus, evaluated with Qwen2.5VL, as shown in Table 5. The overall performance remains at comparable level.
Top-$k$ RAG iNat21 iNat24
k=1 41.98 42.72
k=3 44.39 42.18
k=5 46.93 45.96
k=7 45.99 46.23
k=10 48.26 45.28
k=15 49.73 48.47
k=20 50.53 48.79
Table 5: Top-$k$ RAG performance using iNat24 images as knowledge corpus, evaluated with Qwen2.5VL.
We also note a related line of research (Tanaka et al. 2023; Van Landeghem et al. 2023) on extracting knowledge from document images (e.g., PDF files, Powerpoint slides), conventionally referred to as OCR-based VQA. Although these tasks also require models to retrieve relevant page and interpret images to answer textual queries, the underlying knowledge is primarily textual—obtained via OCR. This focus diverges considerably from our goal of visual feature extraction, where visual cues are inherently more sparse in the wild.
Appendix B Additional Ablation Studies
B.1 Can the answers for Visual-RAG be found in Wikipedia?
We construct a small textual knowledge base from Wikipedia articles for the 10k species in iNat21 dataset. This minimum knowledge base comprising approximately 97,000 chunks, each consisting of $\sim$150 tokens.
@@ B.1 Can the answers for Visual-RAG be found in Wikipedia?
We then applied basic text chunking to these articles, splitting them into chunks of approximately 150 tokens. Each chunk is bounded within a single section of the original article to preserve contextual integrity. Sections such as “Reference”, “See also”, “External links”, etc., which lack informative content were excluded. The chunks are cut upon period mark, so that they do not end or start with incomplete sentence. Two or more adjacent sections having total length less than 150 tokens are merged, with the section titles kept.
For text retrieval, we use a pre-trained dense text retriever, DRAGON Plus (Lin et al. 2023). Given that the ground-truth labels for relevant text chunks are not available, we only assess whether the retrieved text chunks correspond to the correct Wikipedia page, and the retriever achieves Hit@10 of 97%. Next, we present the QA performance across open-sourced models augmented with top-10 retrieved textual knowledge chunks. As shown in Table 6, augmenting with textual knowledge from Wikipedia result in performance drop for both models tested. Instead, in numerous cases, models fail to locate relevant information within the augmented text and output “I don’t know”. This behaviour in turn leads to scores lower than the zero-shot setting.
These experiments provide strong evidence that the knowledge required to answer the questions related to visual features in our benchmark is unlikely to be textually documented in Wikipedia.
Qwen2.5VL GPT-4o
Zero-shot 40.23 56.58
RAG-Best 51.88(@20) 56.58(@7)
Text RAG 25.53 31.68
1GT 41.73 59.89
2GT 48.12 53.38
3GT 54.51 54.89
4GT 53.20 56.77
5GT 52.44 60.15
1-in-10 45.90 56.94
2-in-10 47.20 58.08
3-in-10 51.13 61.09
4-in-10 52.19 59.21
5-in-10 52.07 61.47
Table 6: Ablation experiment: results when augmenting textual knowledge and impact from number of clue images. 1GT - 5GT: input 1-5 ground truth clue images only (mmGT setting); 1-in-10: 1 clue with 9 non-clues, 2 clues with 8 non-clues, etc.
B.2 Does including more clue images consistently lead to better performance?
We observed that in most models, top-$k$ RAG scores at higher $k$ values excel their 1-in-$k$ score. A likely explanation is that the expected number of clue images retrieved in top-$k$ is substantially higher than when only a single clue image is guaranteed. We in turn ablate with effect of multiple clue images. We compare two regimes on the subset of queries with sufficient clues: mmGT ($m$ ground-truth clue images) and $m$-in-10 ($m$ clues mixed with 10m10{-}m non-clues from the same species), reported in Table 6. Aligned with main experiments, mixing non-clues is beneficial at low $m$: non-clue context appears to regularize both models’ multi-image aggregation ability, improving robustness when only a few clues are present. For Qwen, the mmGT results forms a bell-shaped curve, where 3 clue images gives the best performance, while $m$-in-10 gives overall increasing curve. This suggests Qwen benefits most from a small set of clean, diverse clue views, and is sensitive to additional non-clue context. GPT-4o, conversely, gives a U-shaped curve for mmGT. Interestingly, small number of clue images corrupts the model’s correct answer when only one clue is present, and mixing them with non-clues brings the performance close to optimal.
Same Species Different Species
k=3 46.85 46.15
k=5 48.04 47.41
k=7 47.78 46.97
k=10 47.43 47.55
k=15 49.57 47.72
k=20 48.55 47.76
Table 7: Impact from using highly irrelevant distractor in 1-in-k setting, evaluated with Qwen2.5VL.
B.3 Hard negatives can be helpful; how about easy negatives?
We observed that for Qwen2.5VL and some proprietary models, the performance of mixing non-clues with a clue image can sometimes surpass the one clue image upper-bound. As an ablation to the experiment findings, we now mix one clue image with k1k-1 highly irrelevant images from subsets that are not the query species. The results in Table 7 indicate that benefit from hard negatives overall surpass that from easy negatives, especially at higher $k$, partially confirming our hypothesis in last experiment of Section 5.3. As an example, for question “What color are the hindwings of Papaipema inquaesita?”, hard negatives displays the moth not showing the hindwings, while easy negatives are images of plants, mammals, etc. Without the normal image showing only forewings, model could be confused even when presented the hindwings.
B.4 Does visual clue reduce “I don’t know” responses?
During qualitative inspection we observed that, at small $k$, proprietary models often answer with phrases such as “I cannot decide” or “Not enough information.” We conventionally name such outputs as “IDK” responses. A natural hypothesis is that adding clue images should increase model confidence and therefore lower the IDK rate. Table 8 confirms this for the proprietary models and for Qwen: as $k$ grows, their IDK rate falls, mirroring the rise in accuracy. By contrast, Phi, Pixtral, and Llama produce few IDK responses even when their accuracy fluctuates, suggesting that these models tend to commit to an answer rather than abstain. This tendency also explains the seemingly counter-intuitive result that Llama’s accuracy in the Non-clue setting exceeds its zero-shot baseline. Although the added non-clue image lacks the required visual evidence, it could provide the model with organism-specific context, prompting it to draw on parametric knowledge to answer.
Model Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
Baselines
Zero-shot (no image) 0.27 0.00 1.60 0.53 11.76 0.00 0.53 3.74
GT clue (1 image) 0.59 9.89 6.52 0.48 2.84 4.60 2.41 5.83
Non-clue (1 image) 2.14 27.27 31.28 1.87 4.81 76.74 71.12 67.11
Top-K Retrieval-Augmented Generation
k=1 0.80 10.43 13.10 0.80 2.67 58.02 47.86 41.18
3 0.80 8.29 8.29 0.00 1.07 27.54 24.60 20.86
5 0.53 3.74 8.82 0.53 1.34 18.45 16.58 12.57
7 0.80 3.74 7.22 1.07 0.53 15.24 14.17 11.23
10 0.27 2.94 6.42 0.27 1.60 15.24 11.50 8.82
15 2.14 2.94 5.08 0.27 1.34 13.64 9.36 7.22
20 1.87 2.14 3.21 - - 10.70 6.15 8.02
1-in-K Augmented Generation
k=3 1.34 6.42 10.64 0.32 1.02 21.76 8.18 16.36
5 1.02 3.21 9.25 0.27 0.85 15.08 5.77 13.10
7 1.07 2.94 10.64 0.37 1.34 12.41 4.49 10.43
10 1.18 2.40 10.32 0.22 1.17 11.39 3.96 9.46
15 2.41 1.39 8.23 0.32 1.39 9.36 2.46 7.33
20 2.73 1.93 4.97 - - 8.66 1.93 7.65
Table 8: Percentages of answering “I don’t know” across models and settings.
@@ Appendix C Annotation details
For organism species selection in query construction, we aim to ensure that the chosen species are relatively under-represented in textual resources.To achieve that, we filter them based on the length of their Wikipedia summary and description sections. Well-known species typically have multiple sections (e.g., habitats, behaviour, etc.), indicating extensive documentation. We therefore include only those species whose summary and description account for no more than 50% length of the entire Wikipedia article, as an attempt on minimizing reliance on prior knowledge: widely recognized species may have detailed descriptions on visual features across LLM pre-training corpora, even if these details are not explicitly mentioned in Wikipedia.
For the query generation, we mentioned that we have gone through a human annotator originated refined filtering and rewriting process. The prompt template is shown in Table 10, and following shows the example prototype queries generated with the sample prompt:
Figure 4: Example image of Bromus catharticus flower anther. Owned by original uploader at www.inaturalist.org, shared under $C$-BY-NC license.
@@
After query generation and rewriting, and the coarse-level filtering described in Section 4.2, we dispatch the validated queries to the next round of human annotation of image labels. Figure 5 shows the annotation interface for image labelling. [^7] Similar interface is used for query filtering and rewriting annotator to browse, except the image matrix is more dense and displaying only thumbnail sized images. Annotators are required to only label “Y”, to indicate that this image contain the queried feature and is a clue to the query. Leaving the checkbox blank will be considered as negative label. This design aims to save annotators’ effort, as clue images are assumed to be the minority. The rough clue image rate predicted by MLLM in the previous step is also provided (“Y_Rate: 0.003” after the question text) as reference for annotators to estimate the query difficulty. In the meanwhile, annotators provide query answers upon finding a clue image, as shown in Figure 6. Annotators are allowed to skip for a query when finding more than 9 clues after the first 27 images, indicating this query likely fail the clue image threshold. The order of images are randomly shuffled.
Figure 5: Annotation interface for image labelling.
Figure 6: Annotation interface for image label and query answer.
[…]
True 86 1 3
Partial 6 14 2
False 6 17 65
Table 9: Confusion matrix of human evaluation v.s. LLM-as-Judge. We consider “partially correct” as wildcard match, and the macro-$F_{1}$ score is 0.92.
Appendix D Prompt Templates
Table 10, 11, 12 list the prompt templates used through out our experiments.
@@ Appendix D Prompt Templates
A likely related paragraph from Wikipedia is provided for you to better understand the organism. Please avoid question regarding visual features explicitly mentioned in the Wikipedia paragraphs. (e.g., paragraph: “The seeds are colored in red” ->> Don’t ask “What color is the seed”; paragraph: “The plant have symbolic seed pods” (but doesn’t mention its color) ->> You may ask “What shape/color is the seed pod”)
@@
1. What is the common color of tip of leaf of wall-screw moss (scientific name: Tortula muralis)? (Explanation: The moss can be distinguished from similar mosses by its “air-pointed” leaves.) (Feature: color of leaf tip)
@@
2. What pattern can be observed on the tail feather of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) from dorsal (back) view? (Feature: pattern on tail feather from dorsal view) … More ICL Examples omitted
@@
Wiki Paragraph: Bromus catharticus is a species of brome grass known by the common names rescuegrass, grazing brome, prairie grass, and Schrader’s bromegrass. The specific epithet catharticus is Latin, meaning cathartic. The common name rescuegrass refers to the ability of the grass to provide forage after harsh droughts or severe winters. The grass has a diploid number of 42 …… Omitted due to space. Actual prompt includes whole description passage.
@@
Table 10: Prompt Template for Query Generation. Texts in italic are not part of prompt.
@@
<<image>> <<image>> <<image>>
@@
Please evaluate the answer to a question, score from 0 to 1. The reference answer is provided, and the reference is usually short phrases or a single keyword. If the student answer is containing the keywords or similar expressions (including similar or close color/pattern), without any additional guessed information, it is full correct. Similar or close color/pattern includes but not limited to the following cases: pale or light color can appear to be yellowish/greyish under different light condition; dark colors like dark brown, dark grey, dark purple can appear close to each other, and may appear as black as well; stripe pattern can appear as band or ring, dotted pattern can etc. If the student answer have missed some important part in the reference answer, please assign partial score. The reference answer can be in the form of a Python list, in this case, any one of the list item is correct.
If student answer contain irrelevant information not related to question, mark it with “Redundant”, but it does not affect score if related parts are correct. (e.g. Question: what shape are leaves of XYZ plant, Student Answer: shape xxx, color yyy, color is Redundant answer)
If student answer contain features not listed in reference answer, deduct 0.5 score and mark it with “Likely Hallucination”. (e.g., Reference Answer: black and white. Student Answer: black white, with yellow dots, “Yellow dots” is not mentioned in reference). The reference answer sometimes contains additional information not asked in question, usually enclosed by brackets (), to help verifying hallucinations (e.g.: “Shape is xxx (color is yyy)”). Not mentioning additional information in answer is not considered wrong. For yes/no question, reference may contain explanations on why giving yes/no, but it is not necessary for student to answer the explanation; however, if student explains and differs with reference, it is considered as hallucination. Answering “I don’t know”, “Not enough information” or similar is considered wrong, and please mark it with “No Answer”.
Format Instructions: Separate the remarks with score using “||”, that is, use the syntax of: “Score: score $|$ Likely Hallucination”, “Score: score”, “Score: score $|$ Likely Hallucination $|$ Redundant”, “Score: 0 $|$ No Answer ”. If any explanation on why giving the score is needed, do not start a new line and append after remark with brackets, e.g. “Score: score $|$ Redundant $|$ (Explanation: abc)”.
@@
Reference Answer: black eye-round feather, white stripe above eyes. (sometimes connected to the white forehead)
@@
Question: What is the typical color of the antennae of Harris’s checkerspot butterfly (scientific name: Chlosyne harrisii)?
Reference Answer: alternating black and white band, with yellow on the tip
Student Answer: The antennae of Harris’s checkerspot butterfly are black with orange-tipped clubs.
Score: 0.5 (Explanation: not mentioning black and white)
Student Answer: The typical color of the antennae of Harris’s checkerspot butterfly is black with white spots.
Score: 0.5 $|$ Likely Hallucination (Explanation: not white spot but band. Not mentioning the tip)
Question: Are the leaves of burro-weed (scientific name: Ambrosia dumosa) usually covered in small hairs?
Reference Answer: yes
Student Answer: Yes, the leaves of burro-weed (Ambrosia dumosa) are typically covered in small hairs, giving them a grayish or whitish-green appearance.
Score: 1 $|$ Redundant
… More ICL examples omitted…
@@
Table 12: Prompt Template For Automatic Evaluation. Note that for evaluation, no images is provided, and it becomes pure-text open-ended QA evaluation. Texts in italic are not part of prompt.